Circular No. 7910

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002by IN ESO 139-G34 M. Salvo and B. Schmidt, Research School of Astronomy and Astro-

physics, Australian National University (ANU), and Mt. Stromlo Observatory, report that a spectrum (range 380–920 nm, resolution ∼ 0.4 nm) of SN 2002by (cf. IAUC 7865, 7866), taken with the ANU 2.3-m telescope (+ double-beam spectrograph) at Siding Spring Observatory on May 18.53 UT, shows that this object is a somewhat unusual type-I supernova a few weeks past maximum light. E. Cappellaro, Osservatorio Astronomico di Capodimonte, comments that this spectrum closely resembles that of the type-Ia SN 1997bp ∼ 5 weeks past maximum light. SN 1997bp is a peculiar type-Ia supernova, in that it was polarized near optical maximum (see Wang et al. 1998, Bull. AAS 30, 1325), an uncommon characteristic in type-Ia supernovae that indicates an asphericity in the ejecta. Radio observations of SN 2002by obtained by P. Price, E. Berger, J. Chapman, and J. Harnett at the Australia Telescope Compact Array on May 19 show no source detected at the position of the supernova, down to a 3σ limit of 0.2 mJy. Salvo et al. conclude that SN 2002by is a spectroscopically peculiar type-Ia event, ≈ 2 months past maximum light.

COMET C/2002 C1 (IKEYA-ZHANG) A. Lecacheux, Observatoire de Paris; and N. Biver, European Space

Agency, ESTEC, on behalf of the Odin Solar System Topical Team and of the Odin Team, report: “The Odin submillimetric space telescope observed water lines in comet C/2002 C1 with a high spectral resolution (80 m/s) from Apr. 21.9 to 28.3 UT. The H_{2}O 110-101 line at 556.936 GHz was detected with a peak integrated line intensity of 21.6 K km s^{−1} on Apr. 26.8 inside the 2′ Odin beam. This corresponds to a preliminary water-production rate of 1.7 × 10^{29} molecules/s, taking into account opacity effects. The 110-101 line at 547.676 GHz of the ^{18}O isotopic variety of water was also detected. The average of Apr. 23.7–28.1 observations yields an integrated line intensity of 0.24 ± 0.02 K km s^{−1}. Within model uncertainties, the inferred H_{2}^{18}O production rate is consistent with the terrestrial 16O/^{18}O ratio.”

SUPERNOVA 2002da W. M. Wood-Vasey refines the limiting magnitude for the May 3 images

(cf. IAUC 7905, last line) to be 20.8.

2002 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green